# PromptBench: A Unified Library for Evaluation of Large Language Models

### v2

@@ 1 Introduction
Large language models (LLMs) have been revolutionizing various aspects of human life and society unprecedentedly. Evaluation is of paramount importance to understand the true capabilities of LLMs, mitigate potential risks, and eventually, benefit society further (Eisenstein, 2023; Chang et al., 2023). Recent efforts have evaluated LLMs from diverse aspects (Liang et al., 2022; Zheng et al., 2023; Li et al., 2023b; Huang et al., 2023). Among the findings, one of the most important is that current LLMs are sensitive to prompts (Wang et al., 2023b), vulnerable to adversarial prompt attacks (Zhu et al., 2023b), and exposed to data contamination (Willig et al., 2023; Zhou et al., 2023b; Zhu et al., 2023a), which pose severe security and privacy issues (Wang et al., 2023a; Simmons, 2022). On top of that, there have been various prompt learning algorithms developed based on different evaluation metrics, such as BDPL (Diao et al., 2022), GrIPS (Prasad et al., 2022) and Plum (Pan et al., 2023). Given the increasing popularity of LLMs, it is indispensable to develop a unified codebase to enable easy, fast, and flexible evaluation.
@@
There are existing libraries such as LlamaIndex (Liu, 2022), semantic kernel (Microsoft, 2023), and LangChain (Chase, 2022). However, LlamaIndex and LangChain enhance LLM applications by incorporating databases and various data sources, enabling context-aware functionalities. Semantic Kernel aims to merge AI services with programming languages for versatile AI app development. Eval-harness (Gao et al., 2023) offers a comprehensive framework for evaluating generative language models, but does not support other evaluations such as adversarial prompt attacks, prompt engineering, and dynamic evaluation. Zeno (Zeno, 2023) is an AI evaluation platform supporting interaction and visualization, but it is not easy to customize. LiteLLM (BerriAI, 2023) implements a unified API call for different LLM service prodiders, but it does not support research topics such as robustness, prompt engineering, and new protocols. Detailed comparisons are shown in Appendix A. Consequently, there is an urgent need for a unified library dedicated to the comprehensive evaluation of diverse aspects, particularly for research purposes.
@@
This paper introduces PromptBench, a unified python library for evaluating LLMs from comprehensive dimensions. [^1] The name “PromptBench” is the same as (Zhu et al., 2023b) where we only evaluated the robustness against adversarial prompt attack. We decided to keep the name and heavily extend that project. It consists of a wide range of LLMs and evaluation datasets, covering diverse tasks, evaluation protocols, adversarial prompt attacks, and prompt engineering techniques. As a holistic library, it also supports several analysis tools for interpreting the results. Our library is designed in a modular fashion, allowing researchers to easily build evaluation pipelines for their own projects. We open-source PromptBench with comprehensive documents and tutorials [^2] https://promptbench.readthedocs.io/en/latest/ to support easy, flexible, and collaborative evaluation. We believe PromptBench could enhance our understanding of LLMs’ capabilities and spur new research within the community.
@@ 2.1 Components
Models. PromptBench supports both open-source and proprietary LLMs and it is open to add more. Currently, it supports Flan-T5-large (Chung et al., 2022), Dolly (Databricks, 2023), Vicuna (Chiang et al., 2023), Llama2 series (Touvron et al., 2023b), Cerebras-GPT (Dey et al., 2023), GPT-NEOX (Black et al., 2022), Flan-UL2 (Brain, 2023), phi-1.5 (Li et al., 2023c), PaLM2 (Anil et al., 2023), ChatGPT (OpenAI, 2023a), and GPT-4 (OpenAI, 2023b). PromptBench provides a unified LLMModel interface to allow easy construction and inference of a model with specified max generating tokens and generating temperature. More details of the supported models are shown in Appendix B.1.
Datasets and tasks. Currently, PromptBench consists of $1212$ diverse tasks with $2222$ public datasets and naturally supports more. It supports: sentiment analysis (SST-2 (Socher et al., 2013)), grammar correctness (CoLA (Warstadt et al., 2018)), duplicate sentence detection (QQP (Wang et al., 2017) and MRPC (Dolan and Brockett, 2005)), natural language inference (MNLI (Williams et al., 2018), QNLI (Wang et al., 2019), RTE (Wang et al., 2019), and WNLI (Levesque et al., 2012)), multi-task knowledge (MMLU dataset (Hendrycks et al., 2021)), reading comprehension (SQuAD V2 dataset (Rajpurkar et al., 2018)), translation (UN Multi(Eisele and Chen, 2010), IWSLT 2017 (Cettolo et al., 2017)), math problem-solving (Math (Saxton et al., 2019) and GSM8K(Cobbe et al., 2021)), logical reasoning (Boolean Expressions (bench authors, 2023)), commonsense reasoning (Commensense QA (Talmor et al., 2019), QASC (Khot et al., 2020), NummerSense (Lin et al., 2020), Date (bench authors, 2023) and Object Tracking (bench authors, 2023)), symbolic reasoning (LastLetterConcat (Wei et al., 2023)), algorithm (Valid Parentheses dataset (bench authors, 2023)) Through the unified DatasetLoader interface, it supports easy and customizable dataset loading and processing.
@@
Prompts and prompt engineering. PromptBench offers a suite of $44$ distinct prompt types, and additionally, users have the flexibility to craft custom prompts using the Prompt interface. Task-oriented prompts are structured to clearly delineate the specific task expected of the model, whereas role-oriented prompts position the model in a defined role, such as an expert, advisor, or translator. These prompt categories are adaptable for both zero-shot and few-shot learning contexts, offering diverse application possibilities. Moreover, PromptBench currently includes $66$ prominent prompt engineering methods: few-shot Chain-of-Thought (Wei et al., 2023), zero-shot Chain-of-Thought (Kojima et al., 2022), EmotionPrompt (Li et al., 2023a), Expert Prompting (Xu et al., 2023), Generated Knowledge (Liu et al., 2022), and Least to Most (Zhou et al., 2023a). Our framework is not only equipped for the easy integration of these existing techniques through the prompt_engineering module but is also actively evolving to encompass a broader spectrum of prompt engineering methods, enhancing its adaptability in varied evaluation scenarios.
Adversarial prompt attacks. To facilitate the investigation of LLMs’ robustness on prompts, PromptBench integrates $77$ types of adversarial prompt attacks (Zhu et al., 2023b): TextBugger (Li et al., 2019), TextFooler (Jin et al., 2019), BertAttack (Li et al., 2020), DeepWordBug (Gao et al., 2018), Checklist (Ribeiro et al., 2020), StressTest (Naik et al., 2018), and semantics (Zhu et al., 2023b). These attacks can be easily called via the prompt_attack interface. It also supports the usage of curated adversarial prompts to efficiently evaluate the robustness of LLMs.
@@
Different evaluation protocols. By default, PromptBench supports the standard protocol, i.e., the direct inference. PromptBench further supports dynamic (Zhu et al., 2023a) and semantic (Liu et al., 2023) evaluation protocols by dynamically generating testing data. It is open to integrate more new protocols to avoid data contamination.
@@ 2.3 Supported research topics
PromptBench is designed mainly for research purpose, thus it is easy to customize for different topics. As shown in Figure 1(b), it supports different evaluation topics from the research community including benchmarks, scenarios, and protocols. In benchmarks research, it supports standard natural language understanding, natural language generation, and reasoning tasks. It can also be extended to support research on AI agent and interdisciplinary study. In scenario research, it supports adversarial and out-of-distribution evaluation, and can also support other topics such as hallucination and bias by changing the metrics and DatasetLoader interface. In protocol, it naturally supports standard and dynamic evaluation, and can further be verified by including measurement theory.
PromptBench offers three leaderboards to allow easy comparison: adversarial prompt attack, prompt engineering, and dynamic evaluation, as shown in Appendix C. Researchers are welcome to submit new results to our platform. Extensibility is shown in Appendix D that allows convenient extension of the framework.
3 Conclusion
We presented PromptBench, a unified framework for LLMs evaluation. The library is designed in modular fashion to allow users build evaluation pipeline by composing different models, tasks, and prompts. It also facilitates several research directions such as prompt engineering, adversarial prompt attacks, and dynamic evaluation. We treat PromptBench as the first step towards assessing the true capabilities and exploring the boundaries of current LLMs, and believe the benchmark and analysis results from our library could shed lights on designing more robust and human-aligned models. PromptBench is a long-term project and will be actively maintained. We welcome contributions from all potential contributors.
@@ Appendix A Comparison with Related Code Libraries
OpenAI Evals (Contributors, 2023b)
A framework for evaluating LLMs or systems built using LLMs.
model, dataset, prompt, eval
Evaluation pipelines Benchmarks
OpenCompass (Contributors, 2023c)
One-stop platform for large model evaluation, aiming to provide a transparent benchmark for foundation models.
model, dataset
Evaluation pipelines Benchmarks Leaderboards
promptfoo (Contributors, 2023a)
A framework for evaluating prompts and large language models.
model, dataset, prompt, eval
Evaluation pipelines
LM Evaluation Harness (Gao et al., 2023)
A framework for evaluation of autoregressive LMs.
model, dataset, prompt, eval
Evaluation pipelines Benchmarks Leaderboards
HELM (Liang et al., 2022)
Holistic Evaluation of Language Models.
model
Evaluation pipelines Benchmark Leaderboard
PromptBench (Ours)
Research-focused evaluation toolkit.
model, dataset, prompt, prompt engineering, eval
Evaluation pipelines Prompt Engineering Prompt attacks Dynamic evaluation Leaderboard
@@ B.1 Models
In this section, we list the LLMs implemented in PromptBench.
@@
  • phi-1.5 and phi-2 (Li et al., 2023c): phi-1.5 is an LLM with 1.3 billion parameters, builds upon the dataset used for phi-1 with the addition of diverse NLP synthetic texts.
@@
  • ChatGPT (OpenAI, 2023a) and GPT-4 (OpenAI, 2023b): OpenAI’s ChatGPT and GPT-4 are advanced iterations of the GPT series, with ChatGPT tailored for interactive tasks and GPT-4 being the most proficient in the series.
@@
  • Gemini (Google, 2023): Gemini is Google’s new family of LLMs consisting of Gemini Nano, Gemini Pro, and Gemini Ultra. Gemini Pro is released via API service.
@@ B.3 Evaluation protocols
DyVal (Zhu et al., 2023a) is an approach for dynamic evaluation of LLMs by creating complexity-tailored evaluation samples on-the-fly, as opposed to relying on static benchmarks. DyVal synthesized seven distinct reasoning tasks, including: (1) Mathematics, focusing on arithmetic calculations and linear equation solving; (2) Logical Reasoning, involving boolean, deductive, and abductive logic; and (3) Algorithmic Analysis, covering reachability and the maximum sum path problem. MSTemp (Liu et al., 2023) stands for the semantic evalaution protocol which generate out-of-distribution samples by relying on evlauator LLMs and word replacement.
@@ B.4.1 Prompts
Zero shot Task oriented
Determine if the given pair of statements can be considered the same by responding with ’equivalent’ or ’not_equivalent’.
Role oriented
As an instrument for question comparison evaluation, consider the questions and determine if their meaning is the same, responding with ’equivalent’ for similar questions or ’not_equivalent’ for different questions.
Few shot Task oriented
Review the sentence below and identify whether its grammar is ’Acceptable’ or ’Unacceptable’: Here are three examples. Sentence: Our friends won’t buy this analysis, let alone the next one we propose. Answer: acceptable. Sentence: One more pseudo generalization and I’m giving up. Answer: acceptable. Sentence: They drank the pub. Answer: unacceptable.
Role oriented
Functioning as a grammar evaluation tool, analyze the given sentence and decide if it is grammatically correct, responding with ’acceptable’ or ’unacceptable’: Here are three examples. Sentence: Our friends won’t buy this analysis, let alone the next one we propose. Answer: acceptable. Sentence: One more pseudo generalization and I’m giving up. Answer: acceptable. Sentence: They drank the pub. Answer: unacceptable.

### v3

@@ 1 Introduction
Large language models (LLMs) are revolutionizing aspects of human life and society, such as medical diagnostics (McDuff et al., 2023; Thirunavukarasu et al., 2024), and educational tools (Ho et al., 2023). Evaluation is of paramount importance to understand the true capabilities of LLMs, mitigate potential risks, and eventually, benefit society further (Eisenstein, 2023; Chang et al., 2023). Recent efforts have evaluated LLMs from diverse aspects (Hendrycks et al., 2021; Liang et al., 2022; Zheng et al., 2023; Li et al., 2023c; Huang et al., 2023; HuggingFace, 2023). Among the findings, one of the most important is that current LLMs are sensitive to prompts (Wang et al., 2023b), vulnerable to adversarial prompt attacks (Zhu et al., 2023b), and exposed to testset data contamination (Willig et al., 2023; Zhou et al., 2023b; Zhu et al., 2023a), which pose severe security and privacy issues (Wang et al., 2023a; Simmons, 2022). On top of that, there have been various prompt learning algorithms developed based on different evaluation metrics, such as BDPL (Diao et al., 2022), GrIPS (Prasad et al., 2022) and Plum (Pan et al., 2023). Given the increasing popularity of LLMs, it is indispensable to develop a unified codebase to enable easy, fast, and flexible evaluation of large foundation models.
@@
There are existing libraries such as LlamaIndex (Liu, 2022), semantic kernel (Microsoft, 2023), and LangChain (Chase, 2022). LlamaIndex and LangChain enhance LLM applications by incorporating databases and various data sources. Semantic Kernel aims to merge AI services with programming languages for versatile AI app development. Eval-harness (Gao et al., 2023) offers a comprehensive framework for evaluating generative language models. Zeno (Zeno, 2023) is an AI evaluation platform supporting interaction and visualization, but it is not easy to customize. LiteLLM (BerriAI, 2023) implements a unified API call for different LLM service prodiders.
@@
This paper introduces PromptBench, a unified Python library to evaluate LLMs from comprehensive dimensions. It is designed to fill the gaps current libraries have, offering comprehensive support not only for standard model evaluations but also for advanced scenarios including adversarial prompt attacks and dynamic evaluations. Its extensible architecture allows for the incorporation of new evaluation protocols, addressing the limitations found in other tools. The detailed comparisons are shown in Appendix A, highlighting how PromptBench provides a more complete toolkit for the nuanced evaluation of language models, especially in research contexts where adaptability and thoroughness are paramount. It consists of a wide range of LLMs and evaluation datasets, covering diverse tasks, evaluation protocols, adversarial prompt attacks, and prompt engineering techniques. As a holistic library, it also supports several analysis tools for interpreting the results. Our library is designed in a modular fashion, allowing researchers to easily build evaluation pipelines for their own projects. We open-source PromptBench with comprehensive documents and tutorials [^1] https://promptbench.readthedocs.io/en/latest/ to support easy, flexible, and collaborative evaluation. We believe PromptBench could enhance our understanding of LLMs and spur new research within the community.
@@ 2.1 Components
Models. PromptBench supports both open-source and proprietary LLMs and VLMs and it is open to add more. Currently, it supports a diverse range of LLMs and VLMs, ranging from Llama2 series (Touvron et al., 2023b), Mixtral series (Jiang et al., 2024), LlaVa series (Liu et al., 2023a) to GPT series (OpenAI, 2023a, b). PromptBench provides unified LLMModel and VLMModel interfaces to allow easy construction and inference of a model with specified max generating tokens and generating temperature. The interfaces also support customized models, including those that have been fine-tuned for specific applications. More details of the supported models are shown in Appendix B.1.
Datasets and tasks. PromptBench comprises a wide array of tasks, currently supporting diverse challenges across $1212$ tasks and $2222$ public datasets, with the capacity for additional expansions. The supported tasks include fundamental NLP tasks such as sentiment analysis, grammar correctness, and duplicate sentence detection, as well as complex challenges involving natural language inference, multi-task knowledge, and reading comprehension. It also covers specialized areas like translation, mathematical problem-solving, and various forms of reasoning—logical, commonsense, symbolic, and algorithmic. For detailed descriptions of each dataset and specific task configurations (see Appendix B.2). The unified DatasetLoader interface facilitates easy and customizable loading and processing of these datasets, enhancing the usability and flexibility of PromptBench.
@@
Prompts and prompt engineering. PromptBench offers a suite of $44$ distinct prompt types, and additionally, users have the flexibility to craft custom prompts using the Prompt interface. Task-oriented prompts are structured to clearly delineate the specific task expected of the model, whereas role-oriented prompts position the model in a defined role, such as an expert, advisor, or translator. These prompt categories are adaptable for both zero-shot and few-shot learning contexts, offering diverse application possibilities. Moreover, PromptBench currently includes $66$ prominent prompt engineering methods, details can be found in Appendix B.4.2. Our framework is not only equipped for the easy integration of these existing techniques through the prompt_engineering module but is also actively evolving to encompass a broader spectrum of prompt engineering methods, enhancing its adaptability in varied evaluation scenarios.
Adversarial prompt attacks. To facilitate the investigation of LLMs’ robustness on prompts, PromptBench integrates $44$ types of attacks: (1) character-level attacks (Li et al., 2019; Gao et al., 2018), which manipulate texts by introducing typos or errors to words; (2) word-level attacks (Jin et al., 2019; Li et al., 2020), which aim to replace words with synonyms or contextually similar words to deceive LLMs; (3) sentence-level attacks (Ribeiro et al., 2020; Naik et al., 2018), which append irrelevant or extraneous sentences to the end of prompts, intending to distract LLMs; (4) semantic-level attacks (Zhu et al., 2023b), which simulate the linguistic behavior of people from different countries. Details can be found in Appendix B.4.3 These attacks can be easily called via the prompt_attack interface. It also supports the usage of curated adversarial prompts to efficiently evaluate robustness.
@@
Different evaluation protocols. By default, PromptBench supports the standard protocol, i.e., the direct inference. PromptBench further supports dynamic (Zhu et al., 2023a) and semantic (Liu et al., 2023b) evaluation protocols by dynamically generating testing data. It is open to integrate more new protocols to avoid data contamination.
@@ 2.3 Supported research topics
Compared to current evaluation libraries, PromptBench is designed mainly for research purpose, thus it is easy to customize for different topics. As shown in Figure 1(b), it supports different evaluation topics from the research community including benchmarks, scenarios, and protocols. In benchmarks research, it supports standard natural language understanding, natural language generation, and reasoning tasks. It can also be extended to support research on AI agent and interdisciplinary study. In scenario research, it supports adversarial and out-of-distribution evaluation, and can also support other topics such as hallucination and bias by changing the metrics and DatasetLoader interface. In protocol, it naturally supports standard and dynamic evaluation, and can further be verified by including measurement theory. PromptBench offers three leaderboards to allow easy comparison: adversarial prompt attack, prompt engineering, and dynamic evaluation, as shown in Appendix C. Researchers are welcome to submit new results to our platform. Extensibility is shown in Appendix D that allows convenient extension of the framework.
3 Conclusion and Discussion
We presented PromptBench, a unified framework for LLMs evaluation, designed in a modular fashion to build evaluation pipelines with various models, tasks, and prompts. It supports research in prompt engineering, adversarial attacks, and dynamic evaluation. PromptBench is the first step in assessing and exploring the capabilities of current LLMs. We believe our benchmark and analysis will inform the design of more robust, human-aligned models. In the future, new datasets, evaluation protocols, prompt variations, and analytical tools will be added into PromptBench. We also welcome any contributions for PromptBench.
Despite its versatility, PromptBench has limitations. It may not cover all evaluation scenarios, and some metrics might miss nuanced performance differences. The framework’s effectiveness depends on the quality and diversity of datasets and prompts. Addressing these limitations is a priority for our ongoing and future work.
@@ Appendix A Comparison with Related Code Libraries
OpenAI Evals (Contributors, 2023b) A framework for evaluating LLMs or systems built using LLMs. model, dataset, prompt, eval Evaluation pipelines Benchmarks
OpenCompass (Contributors, 2023c) One-stop platform for large model evaluation, aiming to provide a transparent benchmark for foundation models. model, dataset Evaluation pipelines Benchmarks Leaderboards
promptfoo (Contributors, 2023a) A framework for evaluating prompts and large language models. model, dataset, prompt, eval Evaluation pipelines
LM Evaluation Harness (Gao et al., 2023) A framework for evaluation of autoregressive LMs. model, dataset, prompt, eval Evaluation pipelines Benchmarks Leaderboards
HELM (Liang et al., 2022) Holistic Evaluation of Language Models. model Evaluation pipelines Benchmark Leaderboard
PromptBench (Ours) Research-focused evaluation toolkit. model, dataset, prompt, prompt engineering, eval Evaluation pipelines Prompt Engineering Prompt attacks Dynamic evaluation Leaderboard
@@ B.1 Models
In this section, we list the LLMs and VLMS implemented in PromptBench.
@@
  • phi-1.5 and phi-2 (Li et al., 2023d): phi-1.5 is an LLM with 1.3 billion parameters, builds upon the dataset used for phi-1 with the addition of diverse NLP synthetic texts.
@@
  • BLIP2 (Li et al., 2023b): This visual-language model is proposed by Junnan Li, Dongxu Li, Silvio Savarese, Steven Hoi. BLIP-2 leverages frozen pre-trained image encoders and large language models (LLMs) by training a lightweight, 12-layer Transformer encoder in between them, achieving excellent performance in various vision-language tasks
  • LLaVA (Liu et al., 2024): LlaVA (Language-Image LLaMA) is a multimodal model combining language and image data. It extends the LLaMA architecture to handle both modalities, enabling tasks like image captioning, visual question answering, and image-based text generation.
  • Qwen-VL series(Bai et al., 2023): Qwen-VL (Qwen Large Vision Language Model) is the multimodal version of the large model series, Qwen (abbr. Tongyi Qianwen), proposed by Alibaba Cloud. Qwen-VL accepts image, text, and bounding box as inputs, outputs text, and bounding box.
  • InternLM-XComposer2-VL (Dong et al., 2024): InternLM-XComposer2 is a cutting-edge vision-language model excelling in free-form text-image composition and comprehension, crafting content from diverse inputs like outlines, detailed specs, and reference images. Using a Partial LoRA (PLoRA) approach, it balances vision understanding and text composition.
@@
  • ChatGPT (OpenAI, 2023a) and GPT-4 (OpenAI, 2023b): OpenAI’s ChatGPT and GPT-4 are advanced iterations of the GPT series. ChatGPT is tailored for interactive tasks, while GPT-4 is the most proficient in the series and supports image input.
@@
  • Gemini (Google, 2023): The Gemini model is a multimodal language model developed by Google AI, capable of extracting insights from a diverse array of data formats, including images, and video.
@@ B.2 Tasks and Datasets
  • VQAv2 (Goyal et al., 2017): Visual Question Answering (VQA) v2.0 is a dataset containing open-ended questions about images. These questions require an understanding of vision, language and commonsense knowledge to answer. It is the second version of the VQA dataset.
  • NoCaps (Agrawal et al., 2019): NoCaps is a benchmark dataset for image captioning models that can describe images containing novel objects from object detection datasets. It consists of 166,100 human-generated captions describing 15,100 images from the Open Images validation and test sets.
  • MMMU (Yue et al., 2023): MMMU is a comprehensive benchmark designed to evaluate multimodal models on massive multi-discipline tasks demanding college-level subject knowledge and deliberate reasoning. MMMU includes 11.5K meticulously collected multimodal questions from college exams, quizzes, and textbooks. These questions span 30 subjects and 183 subfields, comprising 32 highly heterogeneous image types, such as charts, diagrams, maps, tables, music sheets, and chemical structures.
  • MathVista (Lu et al., 2023): MathVista is a comprehensive benchmark for mathematical reasoning in visual contexts. It includes three new datasets: IQTest, FunctionQA, and PaperQA. These datasets cover various visual domains and are designed to evaluate logical reasoning on puzzle test figures, algebraic reasoning using functional plots, and scientific reasoning with academic paper figures, respectively.
  • AI2D (Kembhavi et al., 2016): AI2D is a dataset of illustrative diagrams for research on diagram understanding and associated question answering. It contains 5000 grade-school science diagrams with over 150,000 rich annotations, their ground truth syntactic parses, and more than 15,000 corresponding multiple-choice questions.
  • ChartQA (Masry et al., 2022): ChartQA is a large-scale dataset of complex reasoning questions over charts that involve visual and logical operations, covering 9.6K human-written questions as well as 23.1K questions generated from human-written chart summaries.
  • ScienceQA (Lu et al., 2022): ScienceQA is a large-scale dataset for multimodal reasoning with diverse science topics and annotations of answers, lectures and explanations. It covers 26 topics, 127 categories and 379 skills from natural science, language science, and social science and so on.
@@ B.3 Evaluation protocols
DyVal (Zhu et al., 2023a) is an approach for dynamic evaluation of LLMs by creating complexity-tailored evaluation samples on-the-fly, as opposed to relying on static benchmarks. DyVal synthesized seven distinct reasoning tasks, including: (1) Mathematics, focusing on arithmetic calculations and linear equation solving; (2) Logical Reasoning, involving boolean, deductive, and abductive logic; and (3) Algorithmic Analysis, covering reachability and the maximum sum path problem. MSTemp (Liu et al., 2023b) stands for the semantic evalaution protocol which generate out-of-distribution samples by relying on evlauator LLMs and word replacement.
@@ B.4.1 Prompts
Zero shot Task oriented Determine if the given pair of statements can be considered the same by responding with ’equivalent’ or ’not_equivalent’.
Role oriented As an instrument for question comparison evaluation, consider the questions and determine if their meaning is the same, responding with ’equivalent’ for similar questions or ’not_equivalent’ for different questions.
Few shot Task oriented Review the sentence below and identify whether its grammar is ’Acceptable’ or ’Unacceptable’: Here are three examples. Sentence: Our friends won’t buy this analysis, let alone the next one we propose. Answer: acceptable. Sentence: One more pseudo generalization and I’m giving up. Answer: acceptable. Sentence: They drank the pub. Answer: unacceptable.
Role oriented Functioning as a grammar evaluation tool, analyze the given sentence and decide if it is grammatically correct, responding with ’acceptable’ or ’unacceptable’: Here are three examples. Sentence: Our friends won’t buy this analysis, let alone the next one we propose. Answer: acceptable. Sentence: One more pseudo generalization and I’m giving up. Answer: acceptable. Sentence: They drank the pub. Answer: unacceptable.